OpenStudy (anonymous):

divide and simplify y^3/y^15 is the answer just guessing y^18 or Y^12

myininaya (myininaya):

y^(3-15)=y^(-12)=1/(y^12)

OpenStudy (mathteacher1729):

http://www.mathsisfun.com/algebra/exponent-laws.html Scroll down for the Exponent laws. :)

OpenStudy (anonymous):

ninja'd^ but yes 1/y^12